### v1

@@ 2.2 Time discretization
To discretize the evolving manifold $M(t)$ we use the reference discretization $\widehat{M}_{h}$ and transport it using a suitable ALE map ${\boldsymbol{\Phi}}^{\mathcal{A}}$ [55, 60, 89, 52, 49, 8, 18, 68, 3]. The reference points $\widehat{{\boldsymbol{y}}}_{i}\in\widehat{M}_{h}$ defining the simplices of the reference triangulation $\mathcal{T}_{h}$ are evolved as ${\boldsymbol{y}}_{i}(t)={\boldsymbol{\Phi}}_{t}^{\mathcal{A}}(\widehat{{\boldsymbol{y}}}_{i})$. This allows to define evolved elements $T(t)$ and facets $F(t)$ such that
(17) $\mathcal{T}_{h}(t)=\{T(t)\},\quad\mathcal{F}_{h}(t)=\{F(t)\},\quad M_{h}(t)=\bigcup T(t),$
@@
(18) $V_{h}(t)=\{\mathrm{span}\{\phi_{i}(t)\}:i=1,\ldots,N.\;\phi_{i}\in V_{h}(\widehat{M}_{h})\}.$
In an analogous way we can also define $V_{h0}(t)$. It follows from the definitions that $V_{h}(0)\equiv V_{h}(\widehat{M}_{h})$. Discrete material velocities are defined by
(19) ${\boldsymbol{v}}_{h}(t,\cdot)=\sum_{i=1}^{N}{\boldsymbol{v}}(t,{\boldsymbol{y}}_{i}(t))\,\phi_{i}(t),\quad{\boldsymbol{v}}^{\mathcal{A}}_{h}(t,\cdot)=\sum_{i=1}^{N}{\boldsymbol{v}}^{\mathcal{A}}(t,{\boldsymbol{y}}_{i}(t))\,\phi_{i}(t),$
@@
(20) $\partial^{\bullet}_{h}\phi_{i}=\partial_{t}\bar{\phi_{i}}+{\boldsymbol{v}}_{h}\cdot\nabla\bar{\phi_{i}},\quad\partial^{\mathcal{A}}_{h}\phi_{i}=\partial_{t}\bar{\phi_{i}}+{\boldsymbol{v}}^{\mathcal{A}}_{h}\cdot\nabla\bar{\phi_{i}}.$
Using an ALE map, which describes the motion of the computational domain, we have that $\partial^{\mathcal{A}}_{h}\phi_{i}=0$.
@@ 3 Structure preservation
In this section we describe how to deal with the challenges highlighted in P.2. In the effort of maintaining the interpretability of the considered systems, we focus on bounds and mass preservation. The method recently proposed in [30, 29] provides a flexible approach to achieve such goals for a wide variety of finite element formulations on stationary bulk domains. Here, we extend those advances to moving surfaces and moving bulk domains. The main advantages of the employed method are:
@@
  • The approach can easily be incorporated as a post-processing step without requiring explicit time integration. While explicit time integration as reviewed in [103] can be effective for hyperbolic equations, here we are also interested in parabolic equations.
  • The simplicity of the algorithm makes it suitable for implementation in legacy codes with negligible overhead computational time with respect to the original non-preserving discretization method. Nevertheless, the method enjoys good stability properties and even allows for error analysis for some specific cases, see [29].
@@
Following the presentation in [29], the derivation of the method starts from a nonlinear PDE in the form
(21) $u_{t}+\mathcal{L}u+\mathcal{N}(u)=0,$
@@
(22) $\partial_{t}u_{h}+\mathcal{L}_{h}u_{h}+\mathcal{N}_{h}(u_{h})=0,$
@@
(23a) $\partial_{t}u_{h}+\mathcal{L}_{h}u_{h}+\mathcal{N}_{h}(u_{h})=\lambda_{h}g^{\prime}(u_{h}),$
(23b) $\lambda_{h}\geq 0,\quad g(u_{h})\geq 0,\quad\lambda_{h}g(u_{h})=0,$
where the second line represents the usual Karush-Kuhn-Tucker conditions for constrained optimization [62, 70]. A core assumption of the scheme presented in [29] is that (23) is satisfied pointwise in a set of points $\Sigma_{h}$ that can be both mesh points or collocation points for $M_{h}$. We must then have:
(24a) $\partial_{t}u_{h}({\boldsymbol{p}})+\mathcal{L}_{h}u_{h}({\boldsymbol{p}})+\mathcal{N}_{h}(u_{h}({\boldsymbol{p}}))=\lambda_{h}({\boldsymbol{p}})g^{\prime}(u_{h}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
(24b) $\lambda_{h}({\boldsymbol{p}})\geq 0,\quad g(u_{h}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}({\boldsymbol{p}})g(u_{h}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$
It is important to notice that $\Sigma_{h}$ does not include points where essential boundary conditions are applied. The next step is to discretize (24) in time, where we here focus on the case of Backward-Euler (BE or BDF-1) time integration in the interval $\{t^{n},t^{n+1}\}$. Higher order extensions are presented in [29] and can be adapted following similar arguments. The key idea is to apply an operator splitting approach to (23) dividing the problem in two steps.
@@
(25) $\frac{\tilde{u}_{h}^{n+1}({\boldsymbol{p}})-u_{h}^{n}({\boldsymbol{p}})}{\tau}+\mathcal{L}_{h}\tilde{u}_{h}^{n+1}({\boldsymbol{p}})+\mathcal{N}_{h}(\tilde{u}_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$
@@
(26a) $\frac{u_{h}^{n+1}({\boldsymbol{p}})-\tilde{u}_{h}^{n+1}({\boldsymbol{p}})}{\tau}=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
(26b) $\lambda_{h}^{n+1}({\boldsymbol{p}})\geq 0,\quad g(u_{h}^{n+1}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}^{n+1}({\boldsymbol{p}})g(u_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$
If we are only interested in bound preservation, it turns out that the solution to (26) is the cutoff function
(27) $u_{h}^{n+1}=\text{cutoff}[\tilde{u}_{h}^{n+1}]=\begin{cases}\tilde{u}_{h}^{n+1}({\boldsymbol{p}}),&\text{ if }a<\tilde{u}_{h}^{n+1}({\boldsymbol{p}})<b,\\ a,&\text{ if }\tilde{u}_{h}^{n+1}({\boldsymbol{p}})<a,\\ b,&\text{ if }\tilde{u}_{h}^{n+1}({\boldsymbol{p}})>b.\end{cases}$
The above step is unfortunately not mass-preserving even if the solution is. To fix this, the authors in [30, 29] introduce an additional global space-independent Lagrange multiplier $\xi_{h}^{n+1}$. The corrector step (26) is modified to include the mass preservation constraint,
(28a) $\frac{1}{\tau}u_{h}^{n+1}({\boldsymbol{p}})=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}}))+\xi_{h}^{n+1}+\frac{1}{\tau}\tilde{u}_{h}^{n+1}({\boldsymbol{p}}),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
(28b) $\lambda_{h}^{n+1}({\boldsymbol{p}})\geq 0,\quad g(u_{h}^{n+1}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}^{n+1}({\boldsymbol{p}})g(u_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
@@
where $(\cdot\,,\,\cdot)^{h}$ is a suitably chosen discrete inner product. In our case this discrete inner product, that has to be expressed pointwise as $(u\,,\,v)^{h}=\sum_{{\boldsymbol{p}}\in\Sigma_{h}}\omega_{\boldsymbol{p}}u({\boldsymbol{p}})v({\boldsymbol{p}})$, is assumed to be equivalent to the classical $L^{2}$ inner product on $V_{h}$. For this reason, whenever required, we will choose $(\cdot\,,\,\cdot)^{h}$ to be the mass-lumped inner product. Rearranging (28a) as
(29) $\frac{1}{\tau}(u_{h}^{n+1}({\boldsymbol{p}})-(\tilde{u}_{h}^{n+1}({\boldsymbol{p}})+\tau\xi_{h}^{n+1}))=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
@@
(30) $F^{n+1}(\xi_{h}^{n+1})=(\text{cutoff}[\tilde{u}_{h}^{n+1}+\tau\xi_{h}^{n+1}]\,,\,1)^{h}-(u_{h}^{n}\,,\,1)^{h}=0.$
Bounds on the energy norm are provided for a class of non-linear PDEs in [29] together with error analysis of a restricted class of problems.
Remark 3.1.
Since $(F^{n+1})^{\prime}$ might not exist, the authors of the scheme in [30, 29] suggest the use of a secant method to find the solution of $F^{n+1}(\xi)=0$ in (30) using
(31) $\xi_{k+1}=\xi_{k}-\frac{F^{n+1}(\xi_{k})(\xi_{k}-\xi_{k-1})}{F^{n+1}(\xi_{k})-F^{n+1}(\xi_{k-1})},$
@@
The technique introduced is clearly directly applicable to a wide variety of equations. The authors themselves in [30, 29] have proposed experiments for Allen-Cahn, Cahn-Hilliard with variable mobility and Fokker-Planck equations on a fixed bulk domain. Inspired by the biophysical mechanisms driving cell reshaping, we will demonstrate their accuracy when modeling complex advection-diffusion-reaction equations and phase-separation phenomena. The structure preservation properties will allow to simultaneously maintain stability, interpretability and physical complexity.
@@ 4 Tangential grid control and ALE discretization
When dealing with evolving domains in a fitted framework, it is not uncommon to incur deformations that lead to highly distorted meshes. This is especially true when gradient flows are involved, since they only prescribe the velocity and hence the displacement in normal direction. Following 11, we introduce a particular ALE map tailored for gradient flow dynamics that works as follows.
Consider the time interval $\{t_{n},t_{n+1}\}$, and suppose the surface $\Gamma_{h}^{n}$ would be displaced, following the physics of the problem, to an updated surface $\widetilde{\Gamma}_{h}^{n+1}$ at time $t_{n+1}$ which is highly distorted. We define an artificial tangential motion aimed at maintaining the surface shape (and consequently energy) but redistributing the nodes more favorably to reduce mesh distortions. We choose the two-stage algorithm proposed in [36] for the following reasons:
@@
The algorithm presented in [36] is based on the assumption that the starting mesh is regular enough. The idea is then to impose an artificial tangential velocity that requires ${\boldsymbol{\Phi}}^{\mathcal{A}}(t,\cdot):\widehat{\Gamma}\to\Gamma$ to be an harmonic map, i.e.
(32) $\begin{cases}\frac{\partial{\boldsymbol{\Phi}}^{\mathcal{A}}}{\partial t}\cdot({\boldsymbol{n}}_{\Gamma}\circ{\boldsymbol{\Phi}}^{\mathcal{A}})=({\boldsymbol{v}}\cdot{\boldsymbol{n}})\circ{\boldsymbol{\Phi}}^{\mathcal{A}},\\ -\Delta_{\,\widehat{\Gamma}}{\boldsymbol{\Phi}}^{\mathcal{A}}=(\varkappa{\boldsymbol{n}})\circ{\boldsymbol{\Phi}}^{\mathcal{A}}.\end{cases}$
Equivalently, ${\boldsymbol{\Phi}}^{\mathcal{A}}$ is required to minimize the energy $\int_{\,\widehat{\Gamma}}|\nabla_{\widehat{M}}{\boldsymbol{\Phi}}^{\mathcal{A}}(t,\cdot)|^{2}$, under the constraint ${\boldsymbol{v}}\cdot{\boldsymbol{n}}_{\Gamma}=0$ on $\Gamma$, which is imposed using the scalar-valued Lagrange multiplier $\varkappa$. The fully discrete version of the scheme reads: given a surface $\widetilde{\Gamma}_{h}^{n+1}$ evolved from an initial surface $\widehat{\Gamma}_{h}$, find ${\boldsymbol{w}}_{h}^{n+1}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\varkappa_{h}^{n+1}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$ such that
@@
(33a) $\qquad=-(\nabla_{\widehat{\Gamma}_{h}}(\mathbf{Id}_{\widetilde{\Gamma}_{h}^{n+1}}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}})\,,\,\nabla_{\widehat{\Gamma}_{h}}({\boldsymbol{\phi}}_{h}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}}))_{\widehat{\Gamma}_{h}}$
(33b) $-({\boldsymbol{w}}_{h}^{n+1}\cdot{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,\psi_{h})_{\widetilde{\Gamma}_{h}^{n+1}}=0,$
@@
(34) $\Delta{\boldsymbol{e}}_{h}^{n+1}$ $=0,\qquad\text{ in }\widetilde{\Omega}^{n+1}_{h},$
(35) ${\boldsymbol{e}}_{h}^{n+1}$ $={\boldsymbol{w}}_{h}^{n+1},\quad\text{ on }\partial\widetilde{\Omega}^{n+1}_{h},$

### v2

@@ 2.2 Time discretization
The evolving manifold $M(t)$ is discretized using the reference discretization $\widehat{M}_{h}$ and transport it using a suitable ALE map ${\boldsymbol{\Phi}}^{\mathcal{A}}$ Hirt et al. [1974], Hughes et al. [1981], Nobile [2001], Gastaldi [2001], Formaggia and Nobile [2004], Badia and Codina [2006], Bonito et al. [2013], Kovács and Power Guerra [2018], Alphonse et al. [2023]. The reference points $\widehat{{\boldsymbol{y}}}_{i}\in\widehat{M}_{h}$ defining the simplices of the reference triangulation $\mathcal{T}_{h}$ are evolved as ${\boldsymbol{y}}_{i}(t)={\boldsymbol{\Phi}}_{t}^{\mathcal{A}}(\widehat{{\boldsymbol{y}}}_{i})$. This allows to define evolved elements $T(t)$ and facets $F(t)$ such that
$\mathcal{T}_{h}(t)=\{T(t)\},\quad\mathcal{F}_{h}(t)=\{F(t)\},\quad M_{h}(t)=\bigcup T(t),$ (17)
@@
$V_{h}(t)=\{\mathrm{span}\{\phi_{i}(t)\}:i=1,\ldots,N.\;\phi_{i}\in V_{h}(\widehat{M}_{h})\}.$ (18)
The space $V_{h0}(t)$ is defined analogously. It follows from the definitions that $V_{h}(0)\equiv V_{h}(\widehat{M}_{h})$. Discrete material velocities are defined by
${\boldsymbol{v}}_{h}(t,\cdot)=\sum_{i=1}^{N}{\boldsymbol{v}}(t,{\boldsymbol{y}}_{i}(t))\,\phi_{i}(t),\quad{\boldsymbol{v}}^{\mathcal{A}}_{h}(t,\cdot)=\sum_{i=1}^{N}{\boldsymbol{v}}^{\mathcal{A}}(t,{\boldsymbol{y}}_{i}(t))\,\phi_{i}(t),$ (19)
@@
$\partial^{\bullet}_{h}\phi_{i}=\partial_{t}\bar{\phi_{i}}+{\boldsymbol{v}}_{h}\cdot\nabla\bar{\phi_{i}},\quad\partial^{\mathcal{A}}_{h}\phi_{i}=\partial_{t}\bar{\phi_{i}}+{\boldsymbol{v}}^{\mathcal{A}}_{h}\cdot\nabla\bar{\phi_{i}}.$ (20)
Using an ALE map, which describes the motion of the computational domain, it holds that $\partial^{\mathcal{A}}_{h}\phi_{i}=0$.
@@ 3 Structure preservation
In this section we describe how to deal with the second one of the challenges highlighted in subsection 1.1, i.e. positivity, bounds and mass preservation in the interest of interpretability. The method recently proposed in Cheng and Shen [2022b, a] provides a flexible approach to achieve such goal for a wide variety of finite element formulations on stationary bulk domains. Here, we extend those advances to moving surfaces and moving bulk domains. The main advantages of the employed method are:
@@
  • The approach can easily be incorporated as a post-processing step without requiring explicit time integration. While explicit time integration as reviewed in Zhang and Shu [2011] can be effective for hyperbolic equations, here we are also interested in parabolic equations.
  • The simplicity of the algorithm makes it suitable for implementation in legacy codes with negligible overhead computational time with respect to the original non-preserving discretization method. Nevertheless, the method enjoys good stability properties and even allows for error analysis for some specific cases, see Cheng and Shen [2022a].
@@
Following the presentation in Cheng and Shen [2022a], the derivation of the method starts from a nonlinear PDE in the form
$u_{t}+\mathcal{L}u+\mathcal{N}(u)=0,$ (21)
@@
$\partial_{t}u_{h}+\mathcal{L}_{h}u_{h}+\mathcal{N}_{h}(u_{h})=0,$ (22)
@@
$\partial_{t}u_{h}+\mathcal{L}_{h}u_{h}+\mathcal{N}_{h}(u_{h})=\lambda_{h}g^{\prime}(u_{h}),$ (23a)
$\lambda_{h}\geq 0,\quad g(u_{h})\geq 0,\quad\lambda_{h}g(u_{h})=0,$ (23b)
where the second line represents the usual Karush-Kuhn-Tucker conditions for constrained optimization Karush [1939], Kuhn and Tucker [1951]. A core assumption of the scheme presented in Cheng and Shen [2022a] is that (23) is satisfied pointwise in a set of points $\Sigma_{h}$ that can be both mesh points or collocation points for $M_{h}$. It must then hold:
$\partial_{t}u_{h}({\boldsymbol{p}})+\mathcal{L}_{h}u_{h}({\boldsymbol{p}})+\mathcal{N}_{h}(u_{h}({\boldsymbol{p}}))=\lambda_{h}({\boldsymbol{p}})g^{\prime}(u_{h}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$ (24a)
$\lambda_{h}({\boldsymbol{p}})\geq 0,\quad g(u_{h}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}({\boldsymbol{p}})g(u_{h}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$ (24b)
It is important to notice that $\Sigma_{h}$ does not include points where essential boundary conditions are applied. The next step is to discretize (24) in time, where we here focus on the case of Backward-Euler (BE or BDF-1) time integration in the interval $t^{n},t^{n+1}]$. Higher order extensions are presented in Cheng and Shen [2022a] and can be adapted following similar arguments. The key idea is to apply an operator splitting approach to (23) dividing the problem in two steps.
@@
$\frac{\tilde{u}_{h}^{n+1}({\boldsymbol{p}})-u_{h}^{n}({\boldsymbol{p}})}{\tau}+\mathcal{L}_{h}\tilde{u}_{h}^{n+1}({\boldsymbol{p}})+\mathcal{N}_{h}(\tilde{u}_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$ (25)
@@
$\frac{u_{h}^{n+1}({\boldsymbol{p}})-\tilde{u}_{h}^{n+1}({\boldsymbol{p}})}{\tau}=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$ (26a)
$\lambda_{h}^{n+1}({\boldsymbol{p}})\geq 0,\quad g(u_{h}^{n+1}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}^{n+1}({\boldsymbol{p}})g(u_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$ (26b)
If one is only interested in bound preservation, it turns out that the solution to (26) is the cutoff function
$u_{h}^{n+1}=\text{cutoff}[\tilde{u}_{h}^{n+1}]=\begin{cases}\tilde{u}_{h}^{n+1}({\boldsymbol{p}}),&\text{ if }a<\tilde{u}_{h}^{n+1}({\boldsymbol{p}})<b,\\ a,&\text{ if }\tilde{u}_{h}^{n+1}({\boldsymbol{p}})<a,\\ b,&\text{ if }\tilde{u}_{h}^{n+1}({\boldsymbol{p}})>b.\end{cases}$ (27)
The above step is unfortunately not mass-preserving even if the solution is. To fix this, the authors in Cheng and Shen [2022b, a] introduce an additional global space-independent Lagrange multiplier $\xi_{h}^{n+1}$. The corrector step (26) is modified to include the mass preservation constraint,
$\frac{1}{\tau}u_{h}^{n+1}({\boldsymbol{p}})=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}}))+\xi_{h}^{n+1}+\frac{1}{\tau}\tilde{u}_{h}^{n+1}({\boldsymbol{p}}),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$ (28a)
$\lambda_{h}^{n+1}({\boldsymbol{p}})\geq 0,\quad g(u_{h}^{n+1}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}^{n+1}({\boldsymbol{p}})g(u_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h},$ (28b)
@@
where $(\cdot\,,\,\cdot)^{h}$ is a suitably chosen discrete inner product. In our case this discrete inner product, that has to be expressed pointwise as $(u\,,\,v)^{h}=\sum_{{\boldsymbol{p}}\in\Sigma_{h}}\omega_{\boldsymbol{p}}u({\boldsymbol{p}})v({\boldsymbol{p}})$, is assumed to be equivalent to the classical $L^{2}$ inner product on $V_{h}$. For this reason, whenever required, $(\cdot\,,\,\cdot)^{h}$ is chosen to be the mass-lumped inner product. Rearranging (28a) as
$\frac{1}{\tau}(u_{h}^{n+1}({\boldsymbol{p}})-(\tilde{u}_{h}^{n+1}({\boldsymbol{p}})+\tau\xi_{h}^{n+1}))=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$ (29)
@@
$F^{n+1}(\xi_{h}^{n+1})=(\text{cutoff}[\tilde{u}_{h}^{n+1}+\tau\xi_{h}^{n+1}]\,,\,1)^{h}-(u_{h}^{n}\,,\,1)^{h}=0.$ (30)
Bounds on the energy norm are provided for a class of non-linear PDEs in Cheng and Shen [2022a] together with error analysis of a restricted class of problems.
Remark 3.
Since $(F^{n+1})^{\prime}$ might not exist, the authors of the scheme in Cheng and Shen [2022b, a] suggest the use of a secant method to find the solution of $F^{n+1}(\xi)=0$ in (30) using
$\xi_{k+1}=\xi_{k}-\frac{F^{n+1}(\xi_{k})(\xi_{k}-\xi_{k-1})}{F^{n+1}(\xi_{k})-F^{n+1}(\xi_{k-1})},$ (31)
@@
The technique introduced is clearly directly applicable to a wide variety of equations. The authors themselves in Cheng and Shen [2022b, a] have proposed experiments for Allen-Cahn, Cahn-Hilliard with variable mobility and Fokker-Planck equations on a fixed bulk domain. Inspired by the biophysical mechanisms driving cell reshaping, we will demonstrate their accuracy when modeling complex advection-diffusion-reaction equations and phase-separation phenomena. The structure preservation properties will allow to simultaneously maintain stability, interpretability and physical complexity.
@@ 4 Tangential grid control and ALE discretization
When dealing with evolving domains in a fitted framework, it is not uncommon to incur deformations that lead to highly distorted meshes, as mentioned in subsection 1.1. This is especially true when gradient flows are involved, since they only prescribe the velocity and hence the displacement in normal direction. Following (11), we introduce a particular ALE map tailored for gradient flow dynamics that works as follows.
Consider the time interval $t_{n},t_{n+1}]$, and suppose the surface $\Gamma_{h}^{n}$ would be displaced, following the physics of the problem, to an updated surface $\widetilde{\Gamma}_{h}^{n+1}$ at time $t_{n+1}$ which is highly distorted. We define an artificial tangential motion aimed at maintaining the surface shape (and consequently energy) but redistributing the nodes more favorably to reduce mesh distortions. We choose the two-stage algorithm proposed in Duan and Li [2024] for the following reasons:
@@
The algorithm presented in Duan and Li [2024] is based on the assumption that the starting mesh is regular enough. The idea is then to impose an artificial tangential velocity that requires ${\boldsymbol{\Phi}}^{\mathcal{A}}(t,\cdot):\widehat{\Gamma}\to\Gamma$ to be an harmonic map, i.e.
$\begin{cases}\frac{\partial{\boldsymbol{\Phi}}^{\mathcal{A}}}{\partial t}\cdot({\boldsymbol{n}}_{\Gamma}\circ{\boldsymbol{\Phi}}^{\mathcal{A}})=({\boldsymbol{v}}\cdot{\boldsymbol{n}})\circ{\boldsymbol{\Phi}}^{\mathcal{A}},\\ -\Delta_{\,\widehat{\Gamma}}{\boldsymbol{\Phi}}^{\mathcal{A}}=(\varkappa{\boldsymbol{n}})\circ{\boldsymbol{\Phi}}^{\mathcal{A}}.\end{cases}$ (32)
Equivalently, ${\boldsymbol{\Phi}}^{\mathcal{A}}$ is required to minimize the energy $\int_{\,\widehat{\Gamma}}|\nabla_{\widehat{M}}{\boldsymbol{\Phi}}^{\mathcal{A}}(t,\cdot)|^{2}$, under the constraint ${\boldsymbol{v}}\cdot{\boldsymbol{n}}_{\Gamma}=0$ on $\Gamma$, which is imposed using the scalar-valued Lagrange multiplier $\varkappa$. The fully discrete version of the scheme reads: given a surface $\widetilde{\Gamma}_{h}^{n+1}$ evolved from an initial reference surface $\widehat{\Gamma}_{h}$, find ${\boldsymbol{w}}_{h}^{n+1}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\varkappa_{h}^{n+1}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$ such that
@@
$\qquad=-(\nabla_{\widehat{\Gamma}_{h}}(\mathbf{Id}_{\widetilde{\Gamma}_{h}^{n+1}}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}})\,,\,\nabla_{\widehat{\Gamma}_{h}}({\boldsymbol{\phi}}_{h}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}}))_{\widehat{\Gamma}_{h}}$ (33a)
$-({\boldsymbol{w}}_{h}^{n+1}\cdot{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,\psi_{h})_{\widetilde{\Gamma}_{h}^{n+1}}=0,$ (33b)
@@
Remark 4.
The algorithm proposed in (33) does not constitute the only option available. Since (33) is a post-processing step, alternative schemes can be used as long as the mesh redistribution is purely tangential Sauer [2014], Hu and Li [2022]. To highlight the flexibility of our framework we further introduce the scheme in Hu and Li [2022], which closely resembles the one presented in Duan and Li [2024]. The fully discrete version of the modified scheme reads: given a surface $\widetilde{\Gamma}_{h}^{n+1}$ evolved from a surface $\Gamma^{n}_{h}$ at the previous timestep, find ${\boldsymbol{w}}_{h}^{n+1}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\varkappa_{h}^{n+1}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$ such that
$\frac{1}{\tau}(\nabla_{\Gamma^{n}_{h}}({\boldsymbol{w}}_{h}^{n+1}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{\tau})\,,\,\nabla_{\Gamma^{n}_{h}}({\boldsymbol{\phi}}_{h}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{\tau}))_{\Gamma^{n}_{h}}-(\varkappa_{h}^{n+1}{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,{\boldsymbol{\phi}}_{h})_{\widetilde{\Gamma}_{h}^{n+1}}$ $=0$ (34a)
$-({\boldsymbol{w}}_{h}^{n+1}\cdot{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,\psi_{h})_{\widetilde{\Gamma}_{h}^{n+1}}$ $=0,$ (34b)
for all ${\boldsymbol{\phi}}_{h}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\psi_{h}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$, where ${\boldsymbol{\Phi}}^{\mathcal{A}}_{\tau}={\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{-t_{n}}$.
@@
$\Delta{\boldsymbol{e}}_{h}^{n+1}$ $=0,\qquad\text{ in }\widetilde{\Omega}^{n+1}_{h},$ (35)
${\boldsymbol{e}}_{h}^{n+1}$ $={\boldsymbol{w}}_{h}^{n+1},\quad\text{ on }\partial\widetilde{\Omega}^{n+1}_{h},$ (36)